### v1

# Improved Recursive QAOA for Solving MAX-CUT on Bipartite Graphs

Abstract
Quantum Approximate Optimization Algorithm (QAOA) is a quantum-classical hybrid algorithm proposed with the goal of approximately solving combinatorial optimization problems such as the MAX-CUT problem. It has been considered a potential candidate for achieving quantum advantage in the Noisy Intermediate-Scale Quantum (NISQ) era and has been extensively studied. However, the performance limitations of low-level QAOA have also been demonstrated across various instances. In this work, we first analytically prove the performance limitations of level-1 QAOA in solving the MAX-CUT problem on bipartite graphs. We derive an upper bound for the approximation ratio based on the average degree of bipartite graphs. Second, we show through numerical results that solving the same problem using level-1 Recursive QAOA (RQAOA), which is one of the variants of QAOA that uses QAOA as a subroutine to reduce the graph size recursively, outperforms the original QAOA but still exhibits limitations as the graph size increases. To address these limitations, we propose a modified RQAOA that reduces the parameter regime optimized in the QAOA subroutine. While reducing the optimization space could generally make it more challenging to find the true optimal parameters, interestingly, we prove that this modified RQAOA can find the exact maximum cut for a parity partitioned graph which includes all weighted bipartite graphs.
I Introduction
Variational Quantum Algorithms (VQAs) Cerezo et al. (2021) have emerged as a practical solution for harnessing quantum computation in the Noisy Intermediate-Scale Quantum (NISQ) era Preskill (2018); Moll et al. (2018). NISQ devices are characterized by their relatively small number of qubits, their susceptibility to noise, and a lack of error correction, which makes them impractical for running traditional deep quantum algorithms like Shor’s algorithm Shor (1997) designed for fault-tolerant quantum systems. VQAs were introduced to address these limitations, employing a hybrid quantum-classical framework Peruzzo et al. (2014); Farhi et al. (a); Moll et al. (2018); Cerezo et al. (2021); Tilly et al. (2022). In these algorithms, the quantum part prepares a parameterized quantum state and measures observables, while the classical part optimizes these parameters to either minimize or maximize an objective function, such as an energy eigenvalue or a cost function. VQAs provide a powerful and practical tool for solving various problems in the NISQ era, from quantum chemistry to optimization and quantum machine learning Peruzzo et al. (2014); Farhi et al. (a); Moll et al. (2018); Schuld and Killoran (2019); Cerezo et al. (2021).
One of the most prominent examples of VQAs is Quantum Approximate Optimization Algorithm (QAOA) Farhi et al. (a). QAOA is designed to tackle combinatorial optimization problems, where the goal is to find an approximate solution that is close to the optimal one. The algorithm works by alternating between applying a problem-specific operator and mixing operators in a parameterized quantum circuit. The parameters of the circuit are optimized using a classical optimizer to improve the solution iteratively. QAOA has gained significant attention for its potential to outperform classical algorithms in solving problems like the MAX-CUT problem, where it aims to find a cut that maximizes the number of edges between two partitions of vertices of a graph Farhi et al. (a); Wang et al. (2018); Cerezo et al. (2021); Zhou et al. (2020). However, despite its promise for near-term quantum applications, it has been established that QAOA, when applied at any constant level, exhibits limited performance in solving the MAX-CUT problem for several instances Hastings ; Bravyi et al. (2020); Farhi et al. (b); Marwaha (2021); Barak and Marwaha . Several variants and extensions of QAOA have been proposed to address its limitations and enhance its performance across a broader range of problems Bravyi et al. (2020); Hadfield et al. (2019); Egger et al. (2021); Herrman et al. (2022); Zhu et al. (2022); Vijendran et al. (2024).
Recursive QAOA (RQAOA), a non-local version of QAOA, is one of the variants of QAOA proposed to enhance the performance on complex instances and scales to larger problems by reducing the problem size Bravyi et al. (2020). The performance of QAOA can be constrained by the $Z_{2}$ symmetry of its quantum states and the geometric locality of the ansatz, which means that the cost operators only involve interactions between nearest neighbor qubits in the underlying graph Bravyi et al. (2020). To address these limitations, RQAOA reduces the problem size iteratively by fixing certain variables based on edge correlation analysis. This iterative reduction can lead to more precise solutions as the recursion progresses through the new connections between previously unlinked nodes, effectively counteracting the locality constraints of QAOA. Although RQAOA has been less extensively investigated compared to other QAOA variants, it is gaining increasing attention as a promising approach for NISQ devices Bravyi et al. (2020, , 2022); Patel et al. (2024); Bae and Lee (2024); Fischer et al. .
It has been shown that RQAOA outperforms the original QAOA for many problem instances. Numerical evidence on random graphs show that even level-1 RQAOA significantly outperforms QAOA at the same level for solving the MAX-CUT problem Bravyi et al. (2020). Another experimental evidence found by the same authors demonstrate that the level-1 RQAOA, which significantly outperforms the level-1 QAOA, is even competitive with Newman’s classical algorithm for general $k$-colorable graphs for solving the MAX-$k$-CUT problem (especially, when $k=3$) which is a generalization of MAX-CUT, where the goal is to partition the vertices of a graph into $k$ disjoint subsets such that the number of edges between different subsets is maximized Bravyi et al. (2022). Furthermore, there have been rigorous proofs supporting this argument that RQAOA surpasses the original QAOA in certain special cases. It has been proved that the level-1 RQAOA achieves the optimal approximation ratio while that of any constant level-$p$ QAOA with $p<n/4$ is less than $1-\frac{1}{4p+2}$ in solving the MAX-CUT problem on cycle graphs, where $n$ is the number of vertices of a cycle graph Bravyi et al. (2020). Recently, a separation of the performance between RQAOA and QAOA for solving the MAX-CUT problem on complete graphs with $2n$ vertices has been proved Bae and Lee (2024). They analytically showed that in this specific case, the level-1 RQAOA achieves an approximation ratio of 1, whereas the approximation ratio of the original QAOA at $p=1$ is strictly upper-bounded as $1-\frac{1}{8n^{2}}$.
Our first significant result in this work supports the claim that RQAOA performs better than QAOA. We have proved the performance limitations of the level-1 QAOA for solving the MAX-CUT problem on bipartite graphs in terms of the approximation ratio. Among the two bounds we derived in Theorem 3, the first bound was shown to be tight for complete bipartite graphs, while the second bound, though slightly looser, was obtained using the average degree which is a key characteristic representing the graph’s density. Our results indicate that, intuitively, the performance of the level-1 QAOA for solving the MAX-CUT problem tends to degrade as a bipartite graph becomes denser and as a complete bipartite graph becomes larger, respectively. In particular, as the size of the complete bipartite graph increases, the performance of the level-1 QAOA becomes worse approaching that of random guessing, in contrast to the case of complete graphs, which converge to one Bae and Lee (2024).
Unlike other provable cases such as cycle graphs Bravyi et al. (2020) and complete graphs Bae and Lee (2024), in bipartite graphs, as the reduction process of RQAOA progresses, the structure of the graph deteriorates, causing it to lose its bipartite nature. This issue becomes more pronounced in graphs with a large number of nodes, where the graph transforms into a weighted graph with each iteration as the weights become more complex. Consequently, analytically proving the performance of RQAOA in such scenarios remains a challenging problem. Although it has not yet been proved whether RQAOA can achieve the optimal performance in solving the MAX-CUT problem on bipartite graphs, we have numerically shown that the level-1 RQAOA has a much better performance than the level-1 QAOA for this problem instance in Sec V.1.
While RQAOA offers crucial improvements over the original QAOA, it also faces several limitations as identified in recent research Patel et al. (2024) and a few works have generalized and extended RQAOA to improve its performance Finžgar et al. (2024); Brady and Hadfield ; Wagner et al. ; Patel et al. (2024). From the numerical results we have obtained for random weighted bipartite graphs with 128 and 256 nodes in Sec V.1, RQAOA appears to fall short of achieving optimal performance. While employing a better optimization method might lead to improved outcomes, it remains unclear whether this issue stems from the limitations of RQAOA itself in determining edge correlations or from challenges in the optimization process during QAOA iterations.
In this paper, we propose a parameter tuning method to enhance the performance of RQAOA and have proven that our modified RQAOA can always find the exact MAX-CUT solution for weighted bipartite graphs. While similar parameter setting schemes applied to QAOA have been proposed Sureshbabu et al. (2024); Shaydulin et al. (2023) and they have focused on finding optimal parameters, our contribution lies in rigorously proving that reducing the parameter search space to preserve the graph structure significantly improves RQAOA performance for certain cases, regardless of whether optimal parameters are found during QAOA iterations.
This paper is organized as follows. In Sec. III, we introduce the MAX-CUT problem, which we aim to solve, and provide a brief overview of the quantum algorithms used to tackle this problem: QAOA and RQAOA. In Sec. IV, we analytically prove that the level-1 QAOA has limited performance in solving the MAX-CUT problem on bipartite graphs, deriving an upper bound for the approximation ratio based on the average degree of the graph. In Sec. V.1, we numerically show that while the level-1 RQAOA outperforms the level-1 QAOA in solving the MAX-CUT problem on bipartite graphs, it also fails to find the exact solution when the graph size increases. We propose a modified RQAOA to improve its performance and prove that our modified RQAOA can exactly solve the MAX-CUT problem on not only weighted bipartite graphs but also more general weighted graphs called a parity partitioned graph $K_{n,m}^{w}$ in Sec. V.2. Finally, we summarize our results and discuss the potential applicability of our modified RQAOA to other instances in Sec. VI.
II Main Results
In this section, we preview our main results. Our first main result is to prove the performance limitations of the level-1 QAOA in solving the MAX-CUT problem on bipartite graphs. The approximation ratio is upper bounded in terms of the average degree which is a key property representing the graph’s density. More details can be found in Sec IV.
Theorem 3. Let $(G,V)$ be a bipartite graph with two disjoint vertex sets $V_{1}$ and $V_{2}$ with $|V_{1}|=n$ and $|V_{2}|=m$ and $E$ be the set of all edges. Let $\alpha$ be the approximation ratio of the level-1 QAOA for solving the MAX-CUT problem on $G$. Then
$\alpha\leq\frac{1}{2}+\sum_{d=1}^{d_{\max}}\frac{d|D_{d}|}{2|E|}\left[\frac{1}{2\sqrt{d}}\left(\frac{\sqrt{d-1}}{\sqrt{d}}\right)^{d-1}\right]\leq\frac{1}{2}+\frac{1}{2\sqrt{e}}\left(\frac{1}{\sqrt{d_{ave}}}+\frac{\sqrt{e}-1}{d_{ave}}\right),$
where $D_{d}$ is the set of vertices with degree $d$ and $d_{ave}=\frac{2|E|}{|V|}$ denotes the average degree of $G$.
In Sec. V.1, we numerically show that the level-1 RQAOA outperforms the level-1 QAOA, and we suspect that RQAOA might not be able to achieve the optimal approximation ratio in solving the MAX-CUT problem on weighted bipartite graphs. Furthermore, it appears that there is an optimization issue in the QAOA subroutine that needs to be addressed for RQAOA to solve the problem with a better approximation ratio. So, we propose a modified RQAOA with a parameter setting strategy to enhance the performance for this problem instances as follows.
Algorithm 2. We only modify the first step of RQAOA1 and the rest part of our algorithm is the same as the original RQAOA1.
  1. 1.
    Apply the level-$1$ QAOA to find the optimal parameters $({\beta}^{*},{\gamma}^{*})$ to maximize $F_{1}(\mathbf{\beta},\mathbf{\gamma})$ in the restricted domain where $0<|\gamma|\leq\frac{\pi}{2w_{ij}^{*}}$ and $0<\beta<\frac{\pi}{4}$ with $w_{ij}^{*}=\max_{(i,j)\in E}|w_{ij}|$.
Our second main result is to rigorously prove that this modified RQAOA can achieve the approximation ratio one in solving the MAX-CUT problem on weighted bipartite graphs.
Theorem 8*. Our modified RQAOA1 can exactly solve the MAX-CUT problem on weighted bipartite graphs.
This theorem is a special case of Theorem 8 and the same result can generealize for a parity partitioned graph $K_{n,m}^{w}$ which includes weighted bipartite graphs as shown in Sec. V.2.
III Premilinaries
III.1 MAX-CUT problem
Let $G=(V,E)$ be a (undirected) graph with the set of vertices $V=\{1,2,\dots,n\}$ and the set of edges ${E=\{(i,j):i,j\in V\}}$. The MAX-CUT problem is a well-known combinatorial optimization problem that aims to split $V$ into two disjoint subsets such that the number of edges spanning the two subsets is maximized. The MAX-CUT problem can be formulated by maximizing the cost function
$C(\mathbf{x})=\frac{1}{2}\sum_{ij\in E}\left(1-x_{i}x_{j}\right)$
for $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})\in\{-1,1\}^{n}$.
III.2 QAOA
QAOA can be viewed as a discrete version of Adiabatic Quantum Computing Farhi et al. (a). The design of QAOA involves constructing a parameterized quantum circuit that alternates between applying the problem Hamiltonian $H_{C}$ (which encodes the optimization problem) and the driving Hamiltonian $H_{B}$ (which ensures broad exploration of the solution space). Here, we only focus on the MAX-CUT problem which can be converted to the following problem Hamiltonian
$H_{C}=\frac{1}{2}\sum_{(i,j)\in E}\left(I-Z_{i}Z_{j}\right),$
where $Z_{i}$ is the Pauli operator $Z$ acting on the $i$-th qubit. The level-$p$ QAOA, denoted by QAOAp, can be described as the following algorithm.
Algorithm 1 (QAOAp Farhi et al. (a)).
The QAOAp is as follows.
  1. 1.
    Prepare the initial state |$\ket{+}^{\otimes n}$.
  2. 2.
    Generate a variational wave function
    |$\ket{\psi_{p}(\mathbf{\beta},\mathbf{\gamma})}=\prod_{j=1}^{p}e^{-i\beta_{j}H_{B}}e^{-i\gamma_{j}H_{C}}\ket{+}^{\otimes n},$
    where $\mathbf{\beta}=(\beta_{1},\ldots,\beta_{p})$, $\mathbf{\gamma}=(\gamma_{1},\ldots,\gamma_{p})$, $H_{C}$ is a problem Hamiltonian, $H_{B}=\sum_{i=1}^{n}X_{i}$ is a driving Hamiltonian, and $X_{i}$ is the Pauli operator $X$ acting on the $i$-th qubit.
  3. 3.
    Compute the expectation value
    $F_{p}(\mathbf{\beta},\mathbf{\gamma})=\bra{\psi_{p}(\mathbf{\beta},\mathbf{\gamma})}H_{C}\ket{\psi_{p}(\mathbf{\beta},\mathbf{\gamma}})=\left<H_{C}\right>$
    by performing the measurement in the computational basis.
  4. 4.
    Apply a classical optimization algorithm to find the optimal parameters
    $(\mathbf{\beta}^{*},\mathbf{\gamma}^{*})=\mathrm{argmax}_{\mathbf{\beta},\mathbf{\gamma}}F_{p}(\mathbf{\beta},\mathbf{\gamma}).$
The approximation ratio $\alpha$ of QAOAp is defined as
$\alpha_{p}=\frac{F_{p}(\mathbf{\beta}^{*},\mathbf{\gamma}^{*})}{C_{\max}},$
where $C_{\max}=\max_{\mathbf{x}\in\{-1,1\}^{n}}C(\mathbf{x})$.
III.3 Recursive QAOA
In this section, we provide a brief overview of RQAOA which was introduced to address the limitations of the original QAOA by incorporating a recursive problem reduction strategy Bravyi et al. (2020). RQAOA iteratively reduces the problem size by fixing specific variables based on edge correlation analysis. As the recursion proceeds, this iterative reduction can yield more accurate solutions by establishing new connections between previously unconnected nodes, thereby counteracting the locality constraints of QAOA. For a level-$p$ RQAOA, denoted by RQAOAp, we consider an Ising-like Hamiltonian, which is closely related the weighted MAX-CUT Hamiltonian, given by
$H_{n}=\sum_{(i,j)\in E}J_{i,j}Z_{i}Z_{j},$
defined on a graph $G_{n}=(V,E)$ with $|V|=n$, where $J_{i,j}\in\mathbb{R}$ are arbitrary. The goal of RQAOA is to approximate
$\max_{\mathbf{x}\in\{-1,1\}^{n}}\bra{\mathbf{x}}H_{n}\ket{\mathbf{x}}.$
Here, |$\ket{\mathbf{x}}=\ket{x_{1},\dots,x_{n}}$ and $Z\ket{x_{i}}=x_{i}\ket{x_{i}}$ for each $i=1,\dots,n$. RQAOAp can be described as the following algorithm.
Algorithm 2 (RQAOAp Bravyi et al. (2020)).
The level-$p$ RQAOA is as follows.
  1. 1.
    Perform the original QAOAp to find the optimal parameters $({\beta}^{*},{\gamma}^{*})$ to maximize $F_{p}(\mathbf{\beta},\mathbf{\gamma})$.
  2. 2.
    Compute the edge expectation values
    $M_{ij}=\bra{\psi_{p}(\mathbf{\beta}^{*},\mathbf{\gamma}^{*})}Z_{i}Z_{j}\ket{\psi_{p}(\mathbf{\beta}^{*},\mathbf{\gamma}^{*})}$
    for every edges $(i,j)\in E$.
  3. 3.
    Pick the edge $(k,l)=\mathrm{argmax}_{(i,j)\in E}M_{ij}$
  4. 4.
    By imposing the constraint $Z_{k}=\textrm{sgn}(M_{kl})Z_{l}$, replace it with $H_{n}$ to obtain
    $H_{n}^{\prime}=\textrm{sgn}(M_{kl})\left[\sum_{(i,k)\in E}J_{ik}Z_{i}Z_{l}\right]+\sum_{i,j\neq k}J_{ij}Z_{i}Z_{j}$
  5. 5.
    Call the QAOA recursively to maximize the expected value of a new Ising Hamiltonian $H_{n-1}$ depending on $n-1$ variables:
    $H_{n-1}=\sum_{(i,l)\in E^{\prime}_{0}}J^{\prime}_{ij}Z_{i}Z_{l}+\sum_{(i,j)\in E^{\prime}_{1}}J^{\prime}_{ij}Z_{i}Z_{j},$
    where
    $E^{\prime}_{0}=\{(i,l):(i,k)\in E\},E^{\prime}_{1}=\{(i,j):(i,j)\neq k\},$
    and
    $J^{\prime}_{ij}=\begin{cases}\mathrm{sgn}(M_{kl})J_{ik}&\mathrm{if}~{}(i,l)\in E^{\prime}_{0},\\ J_{ij}&\mathrm{if}~{}(i,j)\in E^{\prime}_{1}.\end{cases}$
  6. 6.
    The recursion stops when the number of variables reaches some threshold value $n_{c}\ll n$, and find $\mathbf{x}^{*}=\mathrm{argmax}_{\mathbf{x}\in\{-1,1\}^{n_{c}}}\bra{\mathbf{x}}H_{n_{c}}\ket{\mathbf{x}}$ by a classical algorithm.
  7. 7.
    Reconstruct the original (approximate) solution $\tilde{\mathbf{x}}\in\{-1,1\}^{n}$ from $\mathbf{x}^{*}$ using the constraints.
IV Limitation of QAOA1
In this section, we analytically prove that the level-1 QAOA has limited performance in solving the MAX-CUT problem on bipartite graphs. We derive the upper bound of the approximation ratio especially related to the average degree which is the one of important properties of graphs.
The analytical form for the expectation value of the level-$1$ QAOA has been known in Ref. Wang et al. (2018). For all $(i,j)\in E$,
$C_{ij}\right>$ $=$ $\psi_{1}(\beta,\gamma)\right|\frac{1}{2}\left(I-Z_{i}Z_{j}\right)\left|\psi_{1}(\beta,\gamma)\right>$
$=$ $\frac{1}{2}-\frac{1}{4}\sin^{2}(2\beta)\cos^{d_{i}+d_{j}-2(n_{ij}+1)}(1-\cos^{n_{ij}}(2\gamma))+\frac{1}{4}\sin(4\beta)\sin\gamma\left(\cos^{d_{i}-1}\gamma+\cos^{d_{j}-1}\gamma\right),$
where $d_{i}$ is the degree of the vertex $i$ and $n_{ij}$ is the number of triangles with the edge $(i,j)$. Since bipartite graphs are triangle-free, Eq. (IV) becomes simpler. In this case, we can rewrite the expectation value of the MAX-CUT Hamiltonian of the level-1 QAOA in terms of the vertex degrees instead of the edges as follows.
$H_{C}\right>$ $=$ $\sum_{(i,j)\in E}\left[\frac{1}{2}+\frac{1}{4}\sin(4\beta)\sin\gamma\left(\cos^{d_{i}-1}\gamma+\cos^{d_{j}-1}\gamma\right)\right]$
$=$ $\frac{1}{2}|E|+\frac{1}{4}\sin(4\beta)\sin\gamma\sum_{d=1}^{d_{\max}}d|D_{d}|\cos^{d-1}\gamma,$
where $D_{d}$ is the set of vertices with degree $d$. For bipartite graphs, we know the optimal cut is the number of all edges. Thus, the approximation ratio is
$\alpha$ $=$ $\max_{\beta,\gamma}\left<H_{C}\right>/C_{\max}$
$=$ $\max_{\gamma}\left[\frac{1}{2}+\frac{1}{2}\sum_{d=1}^{d_{\max}}\frac{d|D_{d}|}{2|E|}\sin\gamma\cos^{d-1}\gamma\right]$
$=$ $\max_{\gamma}\left[\frac{1}{2}+\frac{1}{2}\sum_{d=1}^{d_{\max}}\frac{d|D_{d}|}{2|E|}f_{d}(\gamma)\right],$
where $f_{d}(\gamma)=\sin\gamma\cos^{d-1}\gamma$. Then we can get a bound of $\alpha$ by obtaining the upper bound of $f_{d}(\gamma)$ for each $d$. Although this bound is obtained quite naturally, but it falls short in capturing the properties of a given bipartite graph. Therefore, by proposing a new bound that utilizes the average degree $d_{ave}$, which reflects one of the important properties of a graph—its density—we can intuitively observe how the density of bipartite graphs impacts the performance of the level-1 QAOA for solving the MAX-CUT problem.
Theorem 3 (Bipartite graph).
Let $(G,V)$ be a bipartite graph with two disjoint vertex sets $V_{1}$ and $V_{2}$ with $|V_{1}|=n$ and $|V_{2}|=m$. Let $\alpha$ be the approximation ratio of the level-1 QAOA for solving the MAX-CUT problem on $G$. Then
$\alpha\leq\frac{1}{2}+\sum_{d=1}^{d_{\max}}\frac{d|D_{d}|}{2|E|}\left[\frac{1}{2\sqrt{d}}\left(\frac{\sqrt{d-1}}{\sqrt{d}}\right)^{d-1}\right]\leq\frac{1}{2}+\frac{1}{2\sqrt{e}}\left(\frac{1}{\sqrt{d_{ave}}}+\frac{\sqrt{e}-1}{d_{ave}}\right),$
where $d_{ave}$ denotes the average vertex degree of $G$ which can be defined as $\frac{\sum_{v\in V}d_{v}}{|V|}$, or equivalently, $\frac{2|E|}{|V|}$.
Proof.
In order to get the upper bound of $\alpha$, let us first find the optimal value of $f_{d}(\gamma)$ as follows. Since $f_{1}(\gamma)=\sin\gamma$, $f_{1}(\gamma)\leq 1$. Observe that $f_{d}(\gamma)=\sin\gamma\cos^{d-1}\gamma$ and $f^{\prime}_{d}(\gamma)=\cos^{d-2}\gamma(d\cos^{2}\gamma-(d-1))$ for each $d\geq 2$. Then we can find the optimal $\gamma^{*}$ such that $\cos\gamma^{*}=(1-\frac{1}{d})^{1/2}$. Thus, we have
$f_{d}(\gamma)\leq f_{d}(\gamma^{*})=\frac{1}{\sqrt{d}}\left(\frac{\sqrt{d-1}}{\sqrt{d}}\right)^{d-1},$
and naturally get the upper bound of the approximation ratio as follows.
$\alpha$$\leq$ $\frac{1}{2}+\sum_{d=1}^{d_{\max}}\frac{d|D_{d}|}{2|E|}\left[\frac{1}{2\sqrt{d}}\left(\frac{\sqrt{d-1}}{\sqrt{d}}\right)^{d-1}\right]$
$\leq$ $\frac{1}{2}+\frac{1}{2}\sum_{d=1}^{d_{\max}}\frac{d|D_{d}|}{2|E|}\left[\sqrt{\frac{1}{ed}}+\frac{1}{d}\left(1-\frac{1}{\sqrt{e}}\right)\right]$
$\leq$ $\frac{1}{2}+\frac{1}{2}\sqrt{\sum_{d=1}^{d_{\max}}\frac{|D_{d}|}{2e|E|}}+\sum_{d=1}^{d_{\max}}\frac{|D_{d}|}{2|E|}\left(1-\frac{1}{\sqrt{e}}\right)$
$\leq$ $\frac{1}{2}+\frac{1}{2\sqrt{e}}\left(\frac{1}{\sqrt{d_{ave}}}+\frac{\sqrt{e}-1}{d_{ave}}\right),$
where $d_{ave}$ denotes the average vertex degree of $G$ which can be defined as $\frac{\sum_{v\in V}d_{v}}{|V|}$, or equivalently, $\frac{2|E|}{|V|}$, or equivalently, $\frac{2|E|}{\sum_{d}|D_{d}|}$. To show that the second inequality holds, we can prove that the following inequalities hold separately
$\frac{1}{\sqrt{d}}\left(\frac{\sqrt{d-1}}{\sqrt{d}}\right)^{d-1}\leq\frac{1}{\sqrt{ed+1-e}}~{}~{}\text{and}~{}~{}\frac{1}{\sqrt{ed+1-e}}\leq\frac{1}{\sqrt{e}d}\left(\sqrt{d}+\sqrt{e}-1\right)$
by showing that
$\ln\left[\frac{1}{\sqrt{d}}\left(\frac{\sqrt{d-1}}{\sqrt{d}}\right)^{d-1}\right]\leq\ln\left(\frac{1}{\sqrt{ed+1-e}}\right)~{}~{}\text{and}~{}~{}ed^{2}\leq(ed+1-e)\left(\sqrt{d}+\sqrt{e}-1\right)^{2}.$
It follows from the concavity of the function
$g(x):=\frac{1}{\sqrt{e}}\sqrt{x}+\left(1-\frac{1}{\sqrt{e}}\right)x$
that the third inequality holds. ∎
Figure 1: Plots illustrating the tightness of the bounds in Theorem 3 and Corollary 4. In (a) and (b), bipartite graphs with $n+m$ vertices with $5\leq n,m\leq 51$ were randomly selected for $p=0.6$ and $p=0.9$, respectively, where $p$ represents the edge probability. The level-$1$ QAOA approximation ratio $\alpha$ was compared with two upper bounds provided in Theorem 3. Compared to (a), the graph instances in (b) are denser, and the bounds are tighter. In (c) and (d), the level-$1$ QAOA approximation ratio $\alpha$ is compared with the bound provided in Corollary 4 for complete bipartite graphs $K_{n,n}$ and $K_{n,m}$, respectively. In this case, it is confirmed that the bounds are tight.
Although the upper bounds we have obtained in Theorem 3 are not tight, they become tighter as a bipartite graph becomes denser which means that $d_{ave}$ increases as you can see in Fig. 1. However, we can get a tight bound for the complete bipartite graphs. Moreover, the intuitive conclusion that can be drawn from the following inequality is that as the size of the complete bipartite graph increases, the performance of the level-1 QAOA deteriorates, approaching that of random guessing.
Corollary 4 (Complete bipartite graph).
Let $K_{n,m}$ be a complete bipartite graph that has two partitioned subsets $V_{1}$ and $V_{2}$ of vertices with $|V_{1}|=n$ and $|V_{2}|=m$.
  • $\bullet$
    For $n=m\geq 2$, $\alpha\leq\frac{1}{2}+\frac{1}{2\sqrt{n-1}}\left(1-\frac{1}{n}\right)^{n/2}$.
  • $\bullet$
    For $n\neq m\geq 2$, $\alpha\leq\frac{1}{2}+\frac{1}{4\sqrt{n-1}}\left(1-\frac{1}{n}\right)^{n/2}+\frac{1}{4\sqrt{m-1}}\left(1-\frac{1}{m}\right)^{m/2}.$
Proof.
For the case of $K_{n,n}$ with $n\geq 2$, all vertices has degree $n$ and so, $|D_{n}|=2n$. Then $\alpha=\max_{\gamma}\left(\frac{1}{2}+\frac{1}{2}f_{n}(\gamma)\right)$. Observe that $f_{n}(\gamma)=2\sin\gamma\cos^{n-1}\gamma$ and $f^{\prime}_{n}(\gamma)=2\cos^{n-2}\gamma(n\cos^{2}\gamma-(n-1))$. Then we can find the optimal $\gamma^{*}$ such that $\cos\gamma^{*}=(1-\frac{1}{n})^{1/2}$ and $f_{n}(\gamma^{*})=\frac{2}{\sqrt{n}}\left(\frac{\sqrt{n-1}}{\sqrt{n}}\right)^{n-1}$. Thus,
$\alpha\leq\frac{1}{2}+\frac{1}{2}f_{n}(\gamma^{*})=\frac{1}{2}+\frac{1}{2\sqrt{n-1}}\left(1-\frac{1}{n}\right)^{n/2}.$
Similarly, for the case of $K_{n,m}$ with $n\neq m\geq 2$, $n$ vertices have degree $m$ and $m$ vertices have degree $n$, that is, $|D_{n}|=m$ and $|D_{m}|=n$. So, $\alpha=\max_{\gamma}\left[\frac{1}{2}+\frac{1}{4}\left(f_{n}(\gamma)+f_{m}(\gamma)\right)\right]$ and thus, we get
$\alpha$$\leq$ $\frac{1}{2}+\frac{1}{4}\left(f_{n}(\gamma^{*})+f_{m}(\gamma^{*})\right)$
$=$ $\frac{1}{2}+\frac{1}{4\sqrt{n-1}}\left(1-\frac{1}{n}\right)^{n/2}+\frac{1}{4\sqrt{m-1}}\left(1-\frac{1}{m}\right)^{m/2}.$
V Improving performance using RQAOA
V.1 RQAOA Performance: Superior to QAOA
In this section, we numerically show a separation of the performance between the level-1 RQAOA and the level-1 QAOA in solving the MAX-CUT problem on weighted bipartite graphs as shown in Fig. 2.
Figure 2: Edge weights were sampled from a Gaussian distribution with a mean of 50 and a variance of 25, subsequently converted to integers. Each benchmark was performed on 10 random instances per $p$. The RQAOA was optimized using a coarse grid search with $20\times 20$ points, followed by gradient descent on the best-found point. The results indicate that RQAOA achieves an approximation ratio of approximately 0.8, while QAOA typically yields an approximation ratio of approximately 0.55 for both graph sizes.
Here, we have selected 10 weighted bipartite graphs with 129 and 256 nodes, where the edge weights were chosen randomly from a Gaussian distribution with a mean of 50 and a variance of 25, subsequently converted to integers. One may think that RQAOA fails to exactly solve the MAX-CUT problem on weighted bipartite graphs. However, there are various factors that could contribute to the observed performance decline of RQAOA. Notably, refining the grid search area more finely is expected to improve RQAOA’s performance. Nonetheless, it remains uncertain whether RQAOA can achieve optimal performance for this problem instance.
Bipartite graphs (both unweighted and weighted) are well-structured and easy to find the exact solution intuitively. Surprisingly, both QAOA and RQAOA have limitations in solving the MAX-CUT problem even for instances like (weighted) bipartite graphs. In the following subsection, we propose a modified RQAOA with a novel parameter-setting strategy to deal with these issues.
V.2 Modified RQAOA achieves the approximation ratio of 1
Normally, if we reduce the region to find the optimal parameters $(\beta,\gamma)$ to maximize the expectation value $F_{p}(\mathbf{\beta},\mathbf{\gamma})$ in QAOA, the maximum expectation value we can obtain from QAOA will be also decreasing. Interestingly, we figured out that in the QAOA subroutine, reducing the optimization domain improves the performance of RQAOA in solving the MAX-CUT problem on weighted bipartite graphs, regardless of whether optimal parameters are found in the QAOA subroutine.
Even though we aim to solve the MAX-CUT problem on unweighted graphs by applying RQAOA, all reduced graphs in each iteration naturally become weighted graphs. Therefore, to analyze the performance, it is crucial to thoroughly understand the performance of the QAOA subroutine on weighted graphs.
In this section, we first define a (undirected) weighted graph with a specific structure $K_{n,m}^{w}$ as follows.
Definition 5.
Let $K_{n,m}^{w}$ be a graph with two partitioned subsets $V_{1}$ and $V_{2}$ of vertices with $|V_{1}|=n$ and $|V_{2}|=m$, and the weight $w_{e}$ of each edge $e$ has the following properties:
  • $\bullet$
    Weight of the edge consisting of vertices belonging to the same vertex set is always negative, that is,
    for the edge $e=(i,j)$, if $i,j\in V_{1}$ or $i,j\in V_{2}$, then $w_{e}\leq 0$.
  • $\bullet$
    Weight of the edge consisting of vertices belonging to the different vertex sets is always positive, that is,
    for the edge $e=(i,j)$, if $(i,j)\in V_{1}\times V_{2}$ , then $w_{e}\geq 0$.
We call $K_{n,m}^{w}$ as a parity partitioned graph. The definition of a parity partitioned graph $K_{n,m}^{w}$ encompasses all weighted bipartite graphs.
Remark 6.
  1. 1.
    If every edge has a non-zero weight, $K_{n,m}^{w}$ must be a weighted complete graph.
  2. 2.
    If every edge in $E_{1}$ and $E_{2}$ has zero weight, $K_{n,m}^{w}$ must be a (complete) weighted bipartite graph.
  3. 3.
    The edge set $E$ of $K_{n,m}^{w}$ can be written as $E=E_{1}\bigcup E_{2}\bigcup E_{3}$, where $E_{1}=V_{1}\times V_{1}$, $E_{2}=V_{2}\times V_{2}$, and $E_{3}=V_{1}\times V_{2}$.
  4. 4.
    The MAX-CUT cost function $C^{w}(x)$ on a weighted graph $G=(V,E,W)$ can be formulated as
    $C^{w}(x)=\frac{1}{2}\sum_{e\in E}w_{e}(1-x_{e}),$
    where $x_{e}=x_{i}x_{j}$ for $e=(i,j)\in E$ with $x_{i}\in\{-1,1\}$ for all $i\in V$, and $w_{e}$ is the edge weight of $e$.
We propose a modified RQAOA equipped with a novel parameter tuning method and prove that it can get the exact solution for solving the MAX-CUT problem on $K_{n,m}^{w}$.
Algorithm 7 (Modified RQAOA1).
We only modify the first step of RQAOA1 and the rest part of our algorithm is the same as the original RQAOA1.
  1. 1.
    Apply the level-$1$ QAOA to find the optimal parameters $({\beta}^{*},{\gamma}^{*})$ to maximize $F_{1}(\mathbf{\beta},\mathbf{\gamma})$ in the restricted domain where $0<|\gamma|\leq\frac{\pi}{2w_{ij}^{*}}$ and $0<\beta<\frac{\pi}{4}$ with $w_{ij}^{*}=\max_{(i,j)\in E}|w_{ij}|$.
Theorem 8.
Our modified RQAOA1 can exactly solve the MAX-CUT problem on a parity partitioned graph $K_{n,m}^{w}$. In particular, it can always get the exact solution of the MAX-CUT problem on weighted bipartite graphs.
We provide the proof of Theorem 8 in the following section.
V.2.1 Proof of Theorem 8
In this section, we prove that applying our parameter tuning method can exactly solve the MAX-CUT problem on a parity partitioned graph $K_{n,m}^{w}$ in the following steps. Lemma 9 states that the MAX-CUT cost function value and the weight structure of $K_{n,m}^{w}$ will be preserved after identifying two connected nodes with the same or opposite sign when they belong to the same or different parts of the vertex set, respectively. Lemma 10 yields that our modified RQAOA assigns edge correlation correctly as described in Lemma 9. Consequently, combining Lemma 9 and Lemma 10, we can prove that our modified RQAOA can always find the exact MAX-CUT solution for a parity partitioned graph $K_{n,m}^{w}$.
Lemma 9.
If we eliminate an edge $e$ by identifying the vertices of $e$ such that the variables of the vertices with the same or opposite sign when $e\in E_{1}\cup E_{2}$ or $e\in E_{3}$, respectively, the value of the MAX-CUT cost function on $K_{n,m}^{w}$ and the weight structure of reduced graphs remain preserved.
We can intuitively see that the weight structure of a parity partitioned graph will be preserved after identifying two connected nodes for both cases as you can see in Fig. 3. Moreover, we can also figure out that the maximum value of the MAX-CUT cost function $C_{n,m}(x)$ will not change. Note that the maximum value of $C_{n,m}(x)$ on $K_{4,3}^{w}$ is the sum of all positive weights, that is, the sum of all red edge weights. In Case 1, the positive weights of the removed edges are maintained by merging with the edge weights of the red bold lines. Thus, the maximum value of $C_{n,m}(x)$ remains to be unchanged. See Appendix A for the detailed proof.
Figure 3: A schematic diagram showing how the MAX-CUT cost function on $K^{w}_{4,3}$ and the weight structure of the graph change after one iteration of our modified RQAOA for the case when we identify (a) two connected nodes in the same partition with the same sign and (b) in the different partitions with different signs; The red/blue edges indicate the edges with the positive/negative weights, respectively. The bold lines represent the edges whose weights change after identifying two vertices. After one iteration, the reduced graph remains to have the same weight structure as shown in the right graph for both cases.
Lemma 10.
If $0<\left|\gamma^{*}\right|\leq\frac{\pi}{2w_{ij}^{*}}$ with $w_{ij}^{*}=\max_{ij\in E}|w_{ij}|$ and $40<\beta^{*}<\pi/$ 4 in the QAOA1 subroutine for solving the MAX-CUT problem on $K_{n,m}^{w}$, the edge expectation values are always positive or negative when the edges belong to $E_{1}\cup E_{2}$ or $E_{3}$, respectively.
Proof.
We can use the analytic form of the edge expectation values in the level-$1$ QAOA subroutine for weighted graphs Vijendran et al. (2024). Let us denote $\mathcal{N}_{i}$ and $\mathcal{T}_{ij}$ as the set of vertices that are adjacent to a vertex $i$ and form a triangle with the edge $(i,j)$, respectively. Then the edge expectation value of the edge $(i,j)$ is
$Z_{i}Z_{j}\right>$ $=$ $-\frac{1}{2}\sin(4\beta)\sin(w_{ij}\gamma)\left(\prod_{k\neq j\in\mathcal{N}_{i}}\cos(w_{ki}\gamma)+\prod_{l\neq i\in\mathcal{N}_{j}}\cos(w_{lj}\gamma)\right)$
$-\frac{1}{2}\sin^{2}(2\beta)\prod_{\begin{subarray}{c}k\neq j\in\mathcal{N}_{i}\\ k\notin\mathcal{T}_{ij}\end{subarray}}\cos(w_{ki}\gamma)\prod_{\begin{subarray}{c}l\neq j\in\mathcal{N}_{j}\\ l\notin\mathcal{T}_{ij}\end{subarray}}\cos(w_{lj}\gamma)\left[\prod_{t\in\mathcal{T}_{ij}}\cos\left((w_{it}+w_{tj})\gamma\right)-\prod_{t\in\mathcal{T}_{ij}}\cos\left((w_{it}-w_{tj})\gamma\right)\right].$
Case 1) Assume that we pick $e^{*}=(i^{*},j^{*})\in E_{1}\cup E_{2}$ to be eliminated. Without loss of generality, let $e^{*}\in E_{1}$. Let $E_{i^{*}}=\{(i^{*},k)\in E:k\neq j^{*}\}$ be the set of edges connected to the vertex $i^{*}$ excluding the edge $e^{*}$, and $F_{i^{*}}=\{(i^{*},k)\in E:k\in\mathcal{T}_{i^{*}j^{*}}\}$ be the set of edges connected to the vertex $i^{*}$, together with the edge $e^{*}$, form triangles. Now, we can rewrite the edge expectation value in terms of the edges with these notations as
$Z_{i^{*}}Z_{j^{*}}\right>$ $=$ $-\frac{1}{2}\sin(4\beta)\sin(w_{e^{*}}^{-}\gamma)\underbrace{\left(\prod_{e_{1}\in E_{i^{*}}\cap E_{1}}\cos(w_{e_{1}}^{-}\gamma)\prod_{e_{3}\in E_{i^{*}}\cap E_{3}}\cos(w_{e_{3}}^{+}\gamma)+\prod_{e_{1}^{\prime}\in E_{j^{*}}\cap E_{1}}\cos(w_{e_{1}^{\prime}}^{-}\gamma)\prod_{e_{3}^{\prime}\in E_{j^{*}}\cap E_{3}}\cos(w_{e_{3}^{\prime}}^{+}\gamma)\right)}_{(*)}$
$-\frac{1}{2}\sin^{2}(2\beta)\underbrace{\prod_{\begin{subarray}{c}e_{1}\in E_{i^{*}}\cap E_{1}\\ e_{1}\notin F_{i^{*}}\end{subarray}}\cos(w_{e_{1}}^{-}\gamma)\prod_{\begin{subarray}{c}e_{3}\in E_{i^{*}}\cap E_{3}\\ e_{3}\notin F_{i^{*}}\end{subarray}}\cos(w_{e_{3}}^{+}\gamma)\prod_{\begin{subarray}{c}e_{1}^{\prime}\in E_{j^{*}}\cap E_{1}\\ e_{1}^{\prime}\notin F_{j^{*}}\end{subarray}}\cos(w_{e_{1}^{\prime}}^{-}\gamma)\prod_{\begin{subarray}{c}e_{3}^{\prime}\in E_{j^{*}}\cap E_{3}\\ e_{3}^{\prime}\notin F_{j^{*}}\end{subarray}}\cos(w_{e_{3}^{\prime}}^{+}\gamma)}_{(**)}$
×$\times\underbrace{\left[\ \prod_{\begin{subarray}{c}f_{1}\in F_{i^{*}}\cap E_{1}\\ f_{3}\in F_{i^{*}}\cap E_{3}\end{subarray}}\cos((w_{f_{1}}^{-}+w_{\tilde{f_{1}}}^{-})\gamma)\cos((w_{f_{3}}^{+}+w_{\tilde{f_{3}}}^{+})\gamma)-\prod_{\begin{subarray}{c}f_{1}\in F_{i^{*}}\cap E_{1}\\ f_{3}\in F_{i^{*}}\cap E_{3}\end{subarray}}\cos((w_{f_{1}}^{-}-w_{\tilde{f_{1}}}^{-})\gamma)\cos((w_{f_{3}}^{+}-w_{\tilde{f_{3}}}^{+})\gamma)\right]}_{(***)},$
where $\tilde{f}$ represents the third edge of the triangle formed by the edges $e^{*}$ and $f$. We denote the negative and positive edge weight of the edge $e$ by $w_{e}^{-}$ and $w_{e}^{+}$, respectively. We can easily see that if $\left|\gamma\right|\leq\frac{\pi}{2w_{ij}^{*}}$ with $w_{ij}^{*}=\max_{ij\in E}|w_{ij}|$, the terms ()(*)( ∗ ) and ()(**)( ∗ ∗ ) are always positive while the term ()(***)( ∗ ∗ ∗ ) is negative because the cosine function is decreasing on that region. Thus, the edge expectation values of the edges belonging to $E_{1}$ (and $E_{2}$) are always positive.
Case 2) Assume that we pick $e^{*}=(i^{*},j^{*})\in E_{3}$ to be eliminated. We can apply a similar argument to the above case. The main differences will be the sign of $\sin(w_{e^{*}}^{+}\gamma)$ and the term ()(***)( ∗ ∗ ∗ ). Let us rewrite the edge expectation value for this case as
$Z_{i^{*}}Z_{j^{*}}\right>$ $=$ $-\frac{1}{2}\sin(4\beta)\sin(w_{e^{*}}^{+}\gamma)\underbrace{\left(\prod_{e_{1}\in E_{i^{*}}\cap E_{1}}\cos(w_{e_{1}}^{-}\gamma)\prod_{e_{3}\in E_{i^{*}}\cap E_{3}}\cos(w_{e_{3}}^{+}\gamma)+\prod_{e_{2}\in E_{j^{*}}\cap E_{2}}\cos(w_{e_{2}}^{-}\gamma)\prod_{e_{3}^{\prime}\in E_{j^{*}}\cap E_{3}}\cos(w_{e_{3}^{\prime}}^{+}\gamma)\right)}_{(*)}$
$-\frac{1}{2}\sin^{2}(2\beta)\underbrace{\prod_{\begin{subarray}{c}e_{1}\in E_{i^{*}}\cap E_{1}\\ e_{1}\notin F_{i^{*}}\end{subarray}}\cos(w_{e_{1}}^{-}\gamma)\prod_{\begin{subarray}{c}e_{3}\in E_{i^{*}}\cap E_{3}\\ e_{3}\notin F_{i^{*}}\end{subarray}}\cos(w_{e_{3}}^{+}\gamma)\prod_{\begin{subarray}{c}e_{2}\in E_{j^{*}}\cap E_{1}\\ e_{2}\notin F_{j^{*}}\end{subarray}}\cos(w_{e_{2}}^{-}\gamma)\prod_{\begin{subarray}{c}e_{3}^{\prime}\in E_{j^{*}}\cap E_{3}\\ e_{3}^{\prime}\notin F_{j^{*}}\end{subarray}}\cos(w_{e_{3}^{\prime}}^{+}\gamma)}_{(**)}$
×$\times\underbrace{\left[\ \prod_{\begin{subarray}{c}f_{1}\in F_{i^{*}}\cap E_{1}\\ f_{3}\in F_{i^{*}}\cap E_{3}\end{subarray}}\cos((w_{f_{1}}^{-}+w_{\tilde{f_{1}}}^{+})\gamma)\cos((w_{f_{3}}^{+}+w_{\tilde{f_{3}}}^{-})\gamma)-\prod_{\begin{subarray}{c}f_{2}\in F_{i^{*}}\cap E_{2}\\ f_{3}\in F_{i^{*}}\cap E_{3}\end{subarray}}\cos((w_{f_{2}}^{-}-w_{\tilde{f_{2}}}^{+})\gamma)\cos((w_{f_{3}}^{+}-w_{\tilde{f_{3}}}^{-})\gamma)\right]}_{(***)}.$
Suppose that $\left|\gamma\right|\leq\frac{\pi}{2w_{ij}^{*}}$ with $w_{ij}^{*}=\max_{ij\in E}w_{ij}$. Similar to Case 1), the terms ()(*)( ∗ ) and ()(**)( ∗ ∗ ) are always positive. However, in contrast to Case 1), $w_{e^{*}}>0$ so that $\sin(w_{e^{*}}\gamma)$ has the opposite sign, and the term ()(***)( ∗ ∗ ∗ ) is positive in this case because the cosine function is decreasing on that region together with the fact that
$|(w_{f_{1}}^{-}+w_{\tilde{f_{1}}}^{+})\gamma|\leq|(w_{f_{2}}^{-}-w_{\tilde{f_{2}}}^{+})\gamma|~{}~{}\text{and}~{}~{}|(w_{f_{3}}^{+}+w_{\tilde{f_{3}}}^{-})\gamma|\leq|(w_{f_{3}}^{+}-w_{\tilde{f_{3}}}^{-})\gamma|.$
Therefore, the edge expectation values of the edges belonging to $E_{3}$ are always negative. ∎
Proof of Theorem 8.
The weighted MAX-CUT cost function on $K_{n,m}^{w}$ can be written as
$C_{n,m}(x)$ $=$ $\frac{1}{2}\sum_{e_{3}\in E_{3}}w_{e_{3}}^{+}+\frac{1}{2}\sum_{e_{1}\in E_{1}}w_{e_{1}}^{-}+\frac{1}{2}\sum_{e_{2}\in E_{2}}w_{e_{2}}^{-}-\frac{1}{2}\sum_{e_{3}\in E_{3}}w_{e_{3}}^{+}x_{e3}-\frac{1}{2}\sum_{e_{1}\in E_{1}}w_{e_{1}}^{-}x_{e_{1}}-\frac{1}{2}\sum_{e_{2}\in E_{2}}w_{e_{2}}^{-}x_{e_{2}},$
where $w_{e}^{+}$ and $w_{e}^{-}$ indicate the positive and negative edge weight of the edge $e$, respectively. It follows from Lemma 9 and Lemma 10 that we can exactly calculate the new MAX-CUT cost function after $k$ iterations. Assume that there are $l$ eliminated edges with the correlated end vertices and $k-l$ eliminated edges with the anti-correlated end vertices during $k$ iterations. Let $\mathcal{X}$ be the subset of $\{-1,1\}^{n+m}$ satisfying those $k$ constraints. Then the maximum value of the new MAX-CUT cost function we can obtain by imposing the $k$ constraints is
$\max_{x\in\mathcal{X}}C_{n,m}(x)$ $=$ $\sum_{e\in E^{\text{elm}}}w_{e}+\max_{x^{\prime}}C_{n-k_{1},m-k_{2}}(x^{\prime})$
$=$ $\sum_{e\in E^{\text{elm}}}w_{e}+\left(\sum_{e_{3}\in E_{3}}w_{e_{3}}^{+}-\sum_{e\in E^{\text{elm}}}w_{e}\right)$
$=$ $\max_{x}C_{n,m}(x),$
where $E^{\text{elm}}$ is the set of all eliminated edge during $k$ iterations and $k=k_{1}+k_{2}$. Thus, by finding the exact MAX-CUT solution of the final reduced graph $K_{n-k_{1},m-k_{2}}^{w}$, the MAX-CUT cost function value of the initial graph will recover perfectly using $k$ constraints. This completes the proof. ∎
VI Conclusion
We have shown the upper bound of the approximation ratio of the level-1 QAOA in solving the MAX-CUT problem on bipartite graphs. Our findings indicate that the performance of the level-1 QAOA generally deteriorates as a bipartite graph becomes denser and as the size of a complete bipartite graph grows.
In addition, we have numerically shown that although RQAOA outperforms QAOA at the same level in solving the MAX-CUT problem on weighted bipartite graphs, there is still room for further performance improvement. To get a better performance of RQAOA, we have proposed a modified RQAOA with a parameter tuning strategy and rigorously prove that it can solve the MAX-CUT problem on a parity partitioned graph $K_{n,m}^{w}$ which contains all weighted bipartite graphs.
There could be several reasons that can explain what aspects of our modified RQAOA lead to its performance improvement. One of them is the limitation of the QAOA subroutine to optimize the parameters. It cannot be guaranteed whether the true optimal parameter is within the reduced region. The recursion step of RQAOA to fix variables on the eliminated edges can be another reason. To analyze the performance of RQAOA more rigorously from this point of view remains to be opened for now.
We figured out that the bad optimization of QAOA would not be the only reason to explain the limitation of RQAOA by showing that RQAOA performs much better even if we reduce the optimization domain for the QAOA subroutine to solve the MAX-CUT problem on weighted bipartite graphs. Especially in the case of well-structured graphs like $K_{n,m}^{w}$, preserving the graph structure across iterations is believed to enhance the RQAOA performance. We have proved that reducing the optimization region yields such effects for these graphs. For future works, we will see if this argument could fit into more general instances such as triangle-free graphs or triangle-free $d$-regular graphs since almost all triangle-free graphs are bipartite Promela et al. (2002).
Appendix A Appendix: Proof of Lemma 9
In this section, we provide the detailed calculations in the proof of Lemma 9 which states that the reduced graph remains to preserve the weight structure of $K_{n,m}^{w}$ and the MAX-CUT cost function will not change after identifying two connected nodes with the same or opposite sign when they belong to the same or different parts of the vertex set, respectively.
To aid the reader’s understanding, the proof is presented under the assumption of a complete parity partitioned graph. However, the same method can also be applied to prove the result for $K_{n,m}^{w}$. Let us denote the cost function of the weighted MAX-CUT problem over $K_{n,m}^{w}$ by
$C_{n,m}(x)$ $=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}$
$-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}},$
where $1_{i}$, $2_{j}$ denote the indices for the $i$th node in $V_{1}$ and the $j$th node in $V_{2}$, and $w_{1_{i}2_{j}}$ denotes the edge weight $w_{e}$ for the edge $e=(v_{1_{j}},v_{2_{j}})$.
Case 1) Assume that we pick $e_{0}\in E_{1}\cup E_{2}$. Without loss of generality, let $e_{0}=(v_{1_{n-1}},v_{1_{n}})\in E_{1}$. We can rewrite the cost function as
$C_{n,m}(x)$ $=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}$
$-\frac{1}{2}\sum_{j=1}^{m}w_{1_{n}2_{j}}\mathbf{x_{1_{n}}}x_{2_{j}}-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i,i^{\prime}\neq n\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}w_{1_{i}1_{n}}x_{1_{i}}\mathbf{x_{1_{n}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}}$
We impose the constraint $x_{1_{n}}=x_{1_{n-1}}$ on the cost function $C_{n,m}(x)$ to get the new cost function
$C_{n,m}^{\prime}(x)$ $=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}$
$-\frac{1}{2}\sum_{j=1}^{m}w_{1_{n}2_{j}}\mathbf{x_{1_{n-1}}}x_{2_{j}}-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}w_{1_{i}1_{n}}x_{1_{i}}\mathbf{x_{1_{n-1}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}}$
$=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}$
$-\frac{1}{2}\sum_{j=1}^{m}w_{1_{n}2_{j}}\mathbf{x_{1_{n-1}}}x_{2_{j}}-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{\mathbf{n-2}}w_{1_{i}1_{n}}x_{1_{i}}\mathbf{x_{1_{n-1}}}-\frac{1}{2}w_{1_{n-1}1_{n}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}}$
$=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}$
$\underbrace{-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}-\frac{1}{2}\sum_{j=1}^{m}w_{1_{n}2_{j}}}_{=\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}}\underbrace{-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{\mathbf{n-2}}w_{1_{i}1_{n}}-\frac{1}{2}w_{1_{n-1}1_{n}}}_{=\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}+C_{n,m}(x^{\prime})$
$=$ $C_{n-1,m}(x^{\prime}).$
Here, $C_{n-1,m}(x^{\prime})$ is the cost function of the reduced complete weighted graph $K_{n-1,m}^{w}$ which can be formulated as
$C_{n-1,m}(x^{\prime})=\frac{1}{2}\sum_{e\in E^{\prime}}w^{\prime}_{e}(1-x^{\prime}_{e}),$
where $E^{\prime}=E-\{e_{0}\}$, $x^{\prime}=(x_{1_{1}},\dots,x_{1_{n-1}},x_{2_{1}},\dots,x_{2_{m}})\in\{-1,1\}^{n+m-1}$, and
$w^{\prime}_{e}=\begin{cases}w_{e}&\mathrm{if}~{}v_{1_{n-1}}\notin e,\\ w_{1_{i}1_{n-1}}+w_{1_{i}1_{n}}&\mathrm{if}~{}e=(v_{1_{i}},v_{1_{n-1}})~{}\mathrm{for}~{}i=1,\cdots,n-2,\\ w_{1_{n-1}2_{j}}+w_{1_{n}2_{j}}&\mathrm{if}~{}e=(v_{1_{n-1}},v_{2_{j}})~{}\mathrm{for}~{}j=1,\cdots,m.\end{cases}$
Case 2) Assume that we pick $e_{0}\in E_{3}$. Without loss of generality, let $e_{0}=(v_{1_{n}},v_{2_{m}})\in E_{3}$. Similarly, we can rewrite the cost function as
$C_{n,m}(x)$ $=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}$
$-\frac{1}{2}\sum_{j=1}^{m}w_{1_{n}2_{j}}\mathbf{x_{1_{n}}}x_{2_{j}}-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i,i^{\prime}\neq n\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}w_{1_{i}1_{n}}x_{1_{i}}\mathbf{x_{1_{n}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}}.$
By imposing the constraint $x_{1_{n}}=-x_{2_{m}}$ on the cost function $C_{n,m}(x)$, we have the new cost function
$C_{n,m}^{\prime}(x)$ $=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}$
$-\frac{1}{2}\sum_{j=1}^{m}(-w_{1_{n}2_{j}})\mathbf{x_{2_{m}}}x_{2_{j}}-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}(-w_{1_{i}1_{n}})x_{1_{i}}\mathbf{x_{2_{m}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}}$
$=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}-\frac{1}{2}\sum_{j=1}^{\mathbf{m-1}}(-w_{1_{n}2_{j}})\mathbf{x_{2_{m}}}x_{2_{j}}$
$-\frac{1}{2}(-w_{1_{n}2_{m}})-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}(-w_{1_{i}1_{n}})x_{1_{i}}\mathbf{x_{2_{m}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}}$
$=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}x_{1_{i}}x_{2_{j}}-\frac{1}{2}\sum_{j=1}^{\mathbf{m-1}}(-w_{1_{n}2_{j}})\mathbf{x_{2_{m}}}x_{2_{j}}$
$-\frac{1}{2}(-w_{1_{n}2_{m}})-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}x_{1_{i}}x_{1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}(-w_{1_{i}1_{n}})x_{1_{i}}\mathbf{x_{2_{m}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}x_{2_{j}}x_{2_{j^{\prime}}}$
$=$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}\underbrace{-\frac{1}{2}\sum_{i=1}^{n-1}\sum_{j=1}^{m}w_{1_{i}2_{j}}-\frac{1}{2}\sum_{j=1}^{\mathbf{m-1}}(-w_{1_{n}2_{j}})-\frac{1}{2}(-w_{1_{n}2_{m}})}_{=-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{1_{i}2_{j}}+\sum_{j=1}^{m}w_{1_{n}2_{j}}}$
$\underbrace{-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n-1}w_{1_{i}1_{i^{\prime}}}-\frac{1}{2}\sum_{i=1}^{n-1}(-w_{1_{i}1_{n}})}_{=-\frac{1}{2}\sum_{\begin{subarray}{c}i,i^{\prime}=1\\ i\neq i^{\prime}\end{subarray}}^{n}w_{1_{i}1_{i^{\prime}}}+\sum_{i=1}^{n-1}w_{1_{i}1_{n}}}-\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}=1\\ j\neq j^{\prime}\end{subarray}}^{m}w_{2_{j}2_{j^{\prime}}}+C_{n-1,m}(x^{\prime})$
$=$ $\sum_{j=1}^{m}w_{1_{n}2_{j}}+\sum_{i=1}^{n-1}w_{1_{i}1_{n}}+C_{n-1,m}(x^{\prime}).$
Here, $C_{n-1,m}(x^{\prime})$ is the cost function of the reduced complete weighted graph $K_{n-1,m}^{w}$ which can be formulated as
$C_{n-1,m}(x^{\prime})=\frac{1}{2}\sum_{e\in E^{\prime}}w^{\prime}_{e}(1-x^{\prime}_{e}),$
where $E^{\prime}=E-\{e_{0}\}$, $x^{\prime}=(x_{1_{1}},\dots,x_{1_{n-1}},x_{2_{1}},\dots,x_{2_{m}})\in\{-1,1\}^{n+m-1}$, and
$w^{\prime}_{e}=\begin{cases}w_{e}&\mathrm{if}~{}v_{1_{n-1}}\notin e,\\ w_{1_{i}2_{m}}-w_{1_{i}1_{n}}&\mathrm{if}~{}e=(v_{1_{i}},v_{2_{m}})~{}\mathrm{for}~{}i=1,\cdots,n-1,\\ -w_{1_{n}2_{j}}+w_{2_{j}2_{m}}&\mathrm{if}~{}e=(v_{2_{j}},v_{2_{m}})~{}\mathrm{for}~{}j=1,\cdots,m-1.\end{cases}$
Note that the sum of edge weights $\sum_{j=1}^{m}w_{1_{n}2_{j}}+\sum_{i=1}^{n-1}w_{1_{i}1_{n}}$ in the new cost function $C^{\prime}_{n,m}(x)$ exactly corresponds to the amount of weight lost when the edge is removed. Therefore, we have also proved that the cost function will not change after identifying two correlated nodes that belong to the different part of the vertex set, and the reduced graph in this case has the same specific weight structure as you can see in Fig. 3.